OpenStudy (anonymous):

Let f(x) = 27x5 – 33x4 – 21x3 and g(x) = 3x2. Find f(x)/g(x) 9x2 – 11x – 7 9x2 + 11x + 7 9x3 – 11x2 – 7x 9x3 + 11x2 + 7x

OpenStudy (agent47):

9x3-11x2-7x

OpenStudy (campbell_st):

this is \[\frac{27x^5 -33x^4 - 21x^3}{3x^2}.... or... \frac{27x^5}{3x^2} - \frac{33x^4}{3x^2} - \frac{21x^3}{3x^2}\] there is a common factor of 3x^2 in each term leaving 9x^3 - 11x^2 - 7x